I have to solve the equation: 12 = t- 6...may I please have some help

Found 2 solutions by THANApHD, praseenakos@yahoo.com:
Answer by THANApHD(104)

12 = t-6
12+6 = t-6+6 (add 6 on both sides)
18= t
t= 18

Answer by praseenakos@yahoo.com(507)

Question:
Solve the equation: 12 = t- 6
Answer:
12 = t- 6

Add 6 on both sides of the expression....

12 + 6 = t - 6 + 6

==> 18 = t


Hence the solution is, t = 18


Hope you understood.


Regards.
Praseena.